OpenStudy (lynfran):

please help me prove this identity im stuck not geting it out at all! (1+tanx+cotx)(cosx-sinx)=(cscx/sec^2x -secx/csc^2x)

OpenStudy (loser66):

just distribute the left hand side, you can get the right hand side.

OpenStudy (lynfran):

im ending up with \[(-\sin^2x/cosx+\cos^2x/sinx)\] when i distribute it

OpenStudy (loser66):

yyyyyyyyyyyyyyyyyyyes

OpenStudy (loser66):

or \(\dfrac{cos^2}{sin}-\dfrac{sin^2}{cos}\) right?

OpenStudy (loser66):

now, \(sin = \dfrac{1}{csc}\) hence \(csc = \dfrac{1}{sin}\) right?

OpenStudy (loser66):

and \(sec (x) =\dfrac{1}{cos(x)} \rightarrow cos(x) =\dfrac{1}{sec(x)}\)

OpenStudy (loser66):

hence the first term is \(\dfrac{csc (x)}{sec^2(x)}\) do the same with the second term. combine to get the right hand side

OpenStudy (loser66):

got it?

OpenStudy (lynfran):

i think so thanks

OpenStudy (loser66):

ok